OpenStudy (anonymous):

Apply the sum and difference formula for sine to simplify the left hand side of the equation sin(x+pi/4) - sin(x-pi/4)=1

OpenStudy (anonymous):

Will medal!!!!

OpenStudy (anonymous):

umm...whut grade r u in?

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

jw

OpenStudy (anonymous):

12th. Senior

OpenStudy (anonymous):

You?

OpenStudy (anonymous):

like this i think sin(A+B) = sin (A)* cos (B) + sin (B) cos (A) sin(A-B) = sin (A) * cos (B) - sin (B) cos (A) therefore, sin(x+pi/4)+sin(x-pi/4) = sin(x)cos(pi/4)+sin(pi/4)cos(x) + sin(x)cos(pi/4)-sin(pi/4)cos(x) = 2*sin(x) cos(pi/4) Now, cos(pi/4)=cos(45 degrees) = 1/sqrt(2). Therefore sin(x+pi/4)+sin(x-pi/4)=2*(1/sqrt(2))*... = sqrt(2) * sin(x) and finally, we make this equal to 1: sin(x+pi/4) +sin(x-pi/4) = sqrt(2) sin(x) = 1 therefore the x's that solve this will be: sin(x)=1/sqrt(2) --> x=pi/4, 3pi/4, 9pi/4, 11pi/4, 15pi/4, 17pi/4, and also -7pi/4, -5pi/4, -13pi/4, -15pi/4...etc in general: x= pi/4 (plus minus) 2*pi and x = - pi/4 (plus minus) 2*pi the "plus minus" 2*pi is considering any number of cycles you want to make around the unit circle.

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

@csheveland does that help

OpenStudy (anonymous):

It does. I am sorry but that Is almost an identical match to the yahoo answer I found. Hmmm.

OpenStudy (anonymous):

thnx!!

OpenStudy (anonymous):

it is the yahoo answer and i am in 9th

OpenStudy (anonymous):

I am graduating at 16 though. So technically I'm a 10th grader. Anyway, thank you for the help.

OpenStudy (anonymous):

I did not mean to sound rude.

OpenStudy (anonymous):

u r very welcome and u didnot sound rude u were asking a question r u n vertual school

OpenStudy (anonymous):

I suck at school doesn't quite fit you......

OpenStudy (anonymous):

haha thx if u wanna talk msg me so i can help my friend ok? thx lol

OpenStudy (anonymous):

and get off of this one

OpenStudy (anonymous):

Sorry. Thnx.

OpenStudy (anonymous):

ur very welcom